SPECTROSCOPIC STUDIES OF THE H_{3}^{+} + H_{2} REACTION AT ASTROPHYSICALLY RELEVANT TEM-PERATURES

BRIAN A. TOM, \underline{BRETT A. McGUIRE,} LAUREN E. MOORE, THOMAS J. WOOD, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801; BENJAMIN J. McCALL, Departments of Chemistry and Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

H_{3}^{+} is the key precursor to ion chemistry in the interstellar medium. It has been employed as an astrophysical probe of conditions of temperature and density due to its ubiquity in a variety of environments. The distribution of ortho- and para- spin modifications of H_{3}^{+} is particularly interesting in this regard. Consequently, it is important to understand the pathways through which changes to the H_{3}^{+} spin distribution can occur. One possible pathway is the H_{3}^{+} + H_{2} → H_{2} + H_{3}^{+} reaction, which proceeds by proton hop and proton exchange and is governed by the conservation of nuclear spin. Cordonnier et al.^{a} studies this reaction at high temperature in a pulsed hollow cathode cell, but to facilitate the understanding of astronomical observations, we need lower temperature measurements. Recently, we have constructed a liquid nitrogen-cooled hollow cathode discharge source and coupled it with multipass absorption spectroscopy to measure the ortho:para ratio of H_{3}^{+} in plasmas at a variety of para-H_{2} enrichment levels at ∼160 K. Previously, we have reported^{b} experimental measurements of the branching ratio between proton hop and exchange in a hydrogenic plasma at ∼80 K. Together, these experiments have allowed us to explore the temperature dependence of this branching ratio and provide valuable information for the interpretation of astronomical observations.

aM. Cordonnier et al., J Chem Phys, 113, 3181 (2000)

bB. A. Tom, M. B. Wiczer, A. A. Mills, K. N. Crabtree, and B. J. McCall, “Observation of nuclear spin selection rules in supersonically expanding plasmas containing H_{3}^{+},” 63rd International Symposium on Molecular Spectroscopy (2008).